# Adaptive Polyak Stepsize with Level-value Adjustment for Distributed Optimization

### 10 Mar 2026 v1

@@ I Introduction
Distributed optimization has gained significant attention as a core computational framework for multi-agent systems, fueled by its extensive use in areas such as smart grids [1, 2, 3], multi-robot networks [4, 5], and federated learning [7, 8]. To improve scalability and communication efficiency, a variety of algorithms have been proposed, ranging from distributed gradient algorithms [9, 10] and mirror descent algorithm [11, 12] to primal-dual algorithms [13, 14]. While these algorithms differ in their design and communication schemes, their practical success remains deeply dependent on a single key element: the selection of an appropriate stepsize policy.
The selection of a suitable stepsize is a core element of optimization algorithms. An excessively large stepsize induces oscillation or even divergence, whereas an overly small one leads to prohibitively slow convergence and poor computational efficiency. Therefore, it is essential to choose a stepsize that allows the algorithm to not only converge but also reach the optimal solution, rather than merely settling within a neighborhood of the optimum. One of the primary challenges in distributed gradient-type algorithms arises from the fact that local gradients evaluated at the global optimum are generally non-zero. To achieve exact convergence to the optimal solution, diminishing stepsize strategies are common, provided they satisfy specific decay conditions [6]. However, these strategies suffer from slow convergence rates, limiting their practical applicability, especially in large-scale networked systems. In contrast, constant stepsizes, while facilitating faster initial convergence, typically guarantee convergence only to a neighborhood of the optimal solution, incurring a persistent steady-state error [6]. To address this fundamental trade-off and ensure exact convergence with constant stepsizes, recent research introduces various correction mechanisms and algorithmic modifications. Prominent among these are techniques such as gradient tracking [15, 16, 17]. Particularly, the practical deployment of these algorithms is often hindered by their heavy reliance on a priori knowledge, such as the connectivity of the network topology and Lipschitz constants. To alleviate this dependency and enhance algorithmic robustness, adaptive strategies receive significant attention, including distributed Barzilai–Borwein algorithms [18, 19] and automated stepsize schemes [20, 21, 22]. Despite these advances, most existing adaptive algorithms still necessitate global information, which is typically unavailable in distributed scenarios. Therefore, developing algorithms capable of efficient and adaptive stepsize adjustment solely based on local information remains a significant and pressing challenge in this field.
@@ III-A Gradient Descent with the Polyak Stepsize Algorithm
where $g_{k}$ denotes the gradient of $f$ at $x_{k}$. If $\|g_{k}\|=0$, then $x_{k+1}=x_{k}$. Intuitively, the core of this stepsize lies in its use of the function-value gap $f(x_{k})-f^{\star}$ as an adaptive scaling factor. This design makes the stepsize automatically adjust according to both the current progress (through $f(x_{k})-f^{\star}$) and the local geometry (through $\|g_{k}\|$). However, the primary challenge of the Polyak stepsize due to the lack of information about the $f^{\star}$. Recently, numerous estimation strategies are proposed to estimate the optimal value $f^{\star}$ in (2), e.g., [32, 33, 34].
@@ III-C DPS-LA Algorithm
$\left(\nabla f_{i}^{k_{i}(j)+t}\right)^{\text{T}}(x-z_{i,k_{i}(j)+t})$
$\leq-\frac{1}{\bar{\gamma}}\beta_{i,k_{i}(j)+t}\|\nabla f_{i}^{k_{i}(j)+t}\|^{2},$ (7)
When local problem (III-C) becomes infeasible, it indicates that agent $i$’s current level-value estimate $\bar{f}_{i}^{k_{i}(j)+\eta}$ is inconsistent with its optimization trajectory. The level-value is then updated to a tighter estimate
@@
6:  if problem (III-C) with added constraint $(\nabla f_{i}(z_{i,k}))^{\text{T}}x\leq(\nabla f_{i}(z_{i,k}))^{\text{T}}z_{i,k}-\frac{1}{\bar{\gamma}}\beta_{i,k}\|\nabla f_{i}(z_{i,k})\|^{2}$ is infeasible then
@@ IV Convergence Analysis
See Appendix -E.
@@
$\min_{M\leq k\leq N}f(\bar{x}_{k})-f(x^{\star})\leq\mathcal{O}\left(\frac{1}{\sqrt{nT}}\right),$
where $M=\lfloor N/2\rfloor$.
@@
See Appendix -F.
@@ V Numerical Experiments
As illustrated in Fig. 4(a), the estimated level-values $\bar{f}_{i}^{k}$ converge rapidly and accurately to the true optimal values $f_{i}(\theta^{\star})$. Furthermore, Fig.4(b) confirms that consensus among agents is also achieved rapidly, contributing to the overall stability and efficiency of the proposed algorithm. This precise estimation enables the algorithm to adaptively compute a more aggressive and near-optimal stepsize, as shown in Fig. 5(a), which significantly accelerates the reduction of the optimality gap. Finally, the results in Fig. 5(b) demonstrate that increasing the number of agents improves the convergence rate, thereby validating the theoretical linear speedup property.
@@ VI Conclusion
In all cases, we have $\alpha_{i,0}\in\left[\frac{\alpha_{0}}{2},\alpha_{0}\right]=\left[\frac{c_{0}\alpha_{0}}{2c_{0}},\frac{c_{0}\alpha_{0}}{c_{0}}\right]$. Hence, the base case holds. Now assume by induction that for some $k$:
@@
Since the constraint system (III-C) admits no feasible solution, $x^{\star}\in\mathcal{X}^{\star}$ is not feasible to (III-C). then the optimal point $x^{\star}\in\mathcal{X}^{\star}$ is necessarily infeasible to (III-C). Consequently, there must exist at least one index $k\in\{k_{i}(j),k_{i}(j)+1,\ldots,k_{i}(j)+\eta\}$ that violates the constraint, such that
@@
-E Proof of Theorem 1
Proof:
$\frac{1}{c_{k}}\leq\frac{\epsilon}{G\sqrt{\sigma\gamma c_{0}\alpha_{0}}}.$
$\alpha_{i,k}\leq\frac{2\epsilon}{3},$
$\frac{1}{n}\sum_{i=1}^{n}\|{\bar{x}_{k}-z_{i,k}}\|\leq\frac{\epsilon}{4G},$
$\left|\|\bar{x}_{k}-x^{\star}\|^{2}-\|x_{i,k}-x^{\star}\|^{2}\right|$
$\leq\frac{\epsilon^{2}}{4},$
from Lemma 5, we analyze the iterate behavior in two cases:
$\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k+1}-x^{*}\|^{2}$
when $k>K_{4}^{\prime}$
Combining (17) and (-E), there exists a final threshold $K_{\text{final}}=\max\{K_{1}^{\prime},K_{2}^{\prime},K_{3}^{\prime},K_{4}^{\prime},\nu\}$, such that for all $k>K_{\text{final}}$, the average squared distance of the iterates to the optimal set converges to zero
-F Proof of Theorem 2
Proof:
$\alpha_{\text{min}}(2-\frac{\gamma}{c_{k}})(f(\bar{x}_{k})-f(x^{\star}))$
$\|z_{i,k}-\bar{x}_{k}\|$
Summing (-F) over $k=M,\dots,T$ and utilizing the telescoping property of the squared distances, we obtain
$\min\limits_{M\leq k\leq T}f(\bar{x}_{k})-f(x^{\star})$
We now analyze the asymptotic behavior of each component in (-F) term by term. For the case $c_{k}=\sqrt{k+1}$, we have
next, we analyze the exponential decay term from $\tau_{k}$
$\sum_{k=M}^{T}\alpha_{\text{min}}(2-\frac{\gamma}{c_{k}})\beta^{k-1}$ (24)
$\sum_{t=0}^{k-2}\beta^{k-t}\frac{1}{\sqrt{t+1}}$
$=\beta^{2}\sum_{t=0}^{k-2}\beta^{k-2-t}\frac{1}{\sqrt{t+1}}$
$=\beta^{2}\sum_{t=0}^{p}\beta^{p-t}\frac{1}{\sqrt{t+1}}$
$\sum_{t=0}^{\lfloor p/2\rfloor}\beta^{p-t}\frac{1}{\sqrt{t+1}}$
$=2\beta^{p/2}(\sqrt{\lfloor p/2\rfloor+1}-1),$
it is easy to verify that for any $1>\beta_{1}>\sqrt{\beta}$

### 20 Mar 2026 v2

@@ I Introduction
Distributed optimization has gained significant attention as a core computational framework for multi-agent systems, fueled by its extensive use in areas such as smart grids [1, 2], multi-robot networks [3]. To improve scalability and communication efficiency, a variety of algorithms have been proposed, ranging from distributed gradient algorithms [4] and primal-dual algorithms [5]. While these algorithms differ in their design and communication schemes, their practical success remains deeply dependent on a single key element: the selection of an appropriate stepsize policy.
The selection of a suitable stepsize is a core element of optimization algorithms. An excessively large stepsize induces oscillation or even divergence, whereas an overly small one leads to prohibitively slow convergence and poor computational efficiency. Therefore, it is essential to choose a stepsize that allows the algorithm to not only converge but also reach the optimal solution, rather than merely settling within a neighborhood of the optimum. One of the primary challenges in distributed gradient-type algorithms arises from the fact that local gradients evaluated at the global optimum are generally non-zero. To achieve exact convergence to the optimal solution, diminishing stepsize strategies are common, provided they satisfy specific decay conditions [6]. However, these strategies suffer from slow convergence rates, limiting their practical applicability, especially in large-scale networked systems. In contrast, constant stepsizes, while facilitating faster initial convergence, typically guarantee convergence only to a neighborhood of the optimal solution, incurring a persistent steady-state error [6]. To address this fundamental trade-off and ensure exact convergence with constant stepsizes, recent research introduces various correction mechanisms and algorithmic modifications. Prominent among these are techniques such as gradient tracking [7, 8]. Particularly, the practical deployment of these algorithms is often hindered by their heavy reliance on a priori knowledge, such as the connectivity of the network topology and Lipschitz constants.
To alleviate this dependency and enhance algorithmic robustness, adaptive strategies receive significant attention, including distributed Barzilai–Borwein algorithms [9, 10] and automated stepsize schemes [11, 12, 13]. Despite these advances, most existing adaptive algorithms still necessitate global information, which is typically unavailable in distributed scenarios. Regarding parameter-free distributed optimization, while [14] focuses on normalized gradient methods, the pioneering work [15] proposes a novel adaptive stepsize mechanism via local backtracking line search. This method operates without any global information and achieves a linear convergence rate. Recent studies [16, 17, 18] further extend this framework to distributed composite optimization with convergence guarantee.
@@ III-A Gradient Descent with the Polyak Stepsize Algorithm
where $g_{k}$ denotes the gradient of $f$ at $x_{k}$. If $\|g_{k}\|=0$, then $x_{k+1}=x_{k}$. Intuitively, the core of this stepsize lies in its use of the function-value gap $f(x_{k})-f^{\star}$ as an adaptive scaling factor. This design makes the stepsize automatically adjust according to both the current progress (through $f(x_{k})-f^{\star}$) and the local geometry (through $\|g_{k}\|$). However, the primary challenge of the Polyak stepsize due to the lack of information about the $f^{\star}$. Recently, numerous estimation strategies are proposed to estimate the optimal value $f^{\star}$ in (2), e.g., [29, 30, 31, 28].
@@ III-C DPS-LA Algorithm
$\left(\nabla f_{i}^{k_{i}(j)+t}\right)^{\text{T}}(x-z_{i,k_{i}(j)+t})\leq-\frac{1}{\bar{\gamma}}\beta_{i,k_{i}(j)+t}\|\nabla f_{i}^{k_{i}(j)+t}\|^{2},$ (7)
When local problem (7) becomes infeasible, it indicates that agent $i$’s current level-value estimate $\bar{f}_{i}^{k_{i}(j)+\eta}$ is inconsistent with its optimization trajectory. The level-value is then updated to a tighter estimate
@@
6:  if problem (7) with added constraint $(\nabla f_{i}(z_{i,k}))^{\text{T}}x\leq(\nabla f_{i}(z_{i,k}))^{\text{T}}z_{i,k}-\frac{1}{\bar{\gamma}}\beta_{i,k}\|\nabla f_{i}(z_{i,k})\|^{2}$ is infeasible then
@@ IV Convergence Analysis
$\frac{1}{c_{k}}\leq\frac{\epsilon}{G\sqrt{\sigma\gamma c_{0}\alpha_{0}}}.$
$\alpha_{i,k}\leq\frac{2\epsilon}{3},$
$\frac{1}{n}\sum_{i=1}^{n}\|{\bar{x}_{k}-z_{i,k}}\|\leq\frac{\epsilon}{4G},$
$\left|\|\bar{x}_{k}-x^{\star}\|^{2}-\|x_{i,k}-x^{\star}\|^{2}\right|$
$\leq\frac{\epsilon^{2}}{4},$
from Lemma 5, we analyze the iterate behavior in two cases:
$\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k+1}-x^{*}\|^{2}$
when $k>K_{4}^{\prime}$
Combining (10) and (IV), there exists a final threshold $K_{\text{final}}=\max\{K_{1}^{\prime},K_{2}^{\prime},K_{3}^{\prime},K_{4}^{\prime},\nu\}$, such that for all $k>K_{\text{final}}$, the average squared distance of the iterates to the optimal set converges to zero
@@
$\min_{M\leq k\leq T}f(\bar{x}_{k})-f(x^{\star})\leq\mathcal{O}\left(\frac{1}{\sqrt{nT}}\right),$
where $M=\lfloor T/2\rfloor$.
@@
$\alpha_{\text{min}}(2-\frac{\gamma}{c_{k}})(f(\bar{x}_{k})-f(x^{\star}))$
$\|z_{i,k}-\bar{x}_{k}\|$
Summing (IV) over $k=M,\dots,T$ and utilizing the telescoping property of the squared distances, we obtain
$\min\limits_{M\leq k\leq T}f(\bar{x}_{k})-f(x^{\star})$
We now analyze the asymptotic behavior of each component in (IV) term by term. For the case $c_{k}=\sqrt{k+1}$, we have
next, we analyze the exponential decay term from $\tau_{k}$
$\sum_{k=M}^{T}\alpha_{\text{min}}(2-\frac{\gamma}{c_{k}})\beta^{k-1}$ (17)
$\sum_{t=0}^{k-2}\beta^{k-t}\frac{1}{\sqrt{t+1}}$
$=\beta^{2}\sum_{t=0}^{k-2}\beta^{k-2-t}\frac{1}{\sqrt{t+1}}$
$=\beta^{2}\sum_{t=0}^{p}\beta^{p-t}\frac{1}{\sqrt{t+1}}$
$\sum_{t=0}^{\lfloor p/2\rfloor}\beta^{p-t}\frac{1}{\sqrt{t+1}}$
$=2\beta^{p/2}(\sqrt{\lfloor p/2\rfloor+1}-1),$
it is easy to verify that for any $1>\beta_{1}>\sqrt{\beta}$
@@ V Numerical Experiments
As illustrated in Fig. 4, the estimated level-values $\bar{f}_{i}^{k}$ converge rapidly and accurately to the true optimal values $f_{i}(\theta^{\star})$. Furthermore, Fig.4(b) confirms that consensus among agents is also achieved rapidly, contributing to the overall stability and efficiency of the proposed algorithm. This precise estimation enables the algorithm to adaptively compute a more aggressive and near-optimal stepsize, as shown in Fig. 5(a), which significantly accelerates the reduction of the optimality gap. Finally, the results in Fig. 5(b) demonstrate that increasing the number of agents improves the convergence rate, thereby validating the theoretical linear speedup property.
@@ VI Conclusion
In all cases, we have $\alpha_{i,0}\in\left[\frac{\alpha_{0}}{2},\alpha_{0}\right]=\left[\frac{c_{0}\alpha_{0}}{2c_{0}},\frac{c_{0}\alpha_{0}}{c_{0}}\right]$. Hence, the base case holds. Now assume by induction that for some $k$,
@@
Since the constraint system (7) admits no feasible solution, $x^{\star}\in\mathcal{X}^{\star}$ is not feasible to (7). then the optimal point $x^{\star}\in\mathcal{X}^{\star}$ is necessarily infeasible to (7). Consequently, there must exist at least one index $k\in\{k_{i}(j),k_{i}(j)+1,\ldots,k_{i}(j)+\eta\}$ that violates the constraint, such that